# Searches for the BSM scenarios at the LHC using decision tree based machine learning algorithms: A comparative study and review of Random Forest, Adaboost, XGboost and LightGBM frameworks

### v1

@@ 3.1.2 Searches for RPV SUSY scenarios using BDT
In the presence of non-zero trilinear couplings $\lambda_{ijk}$, the LSP decays into $l_{k}^{\prime\pm}l_{i/j}^{\prime\mp}\nu_{j/i}$, where $l^{\prime}=e,\mu,\tau$. Thus, the final states become leptonically enriched. For different choices of $\lambda_{ijk}$ couplings, the LHC collaborations have already derived limits on chargino and slepton masses (ATLAS:2021yyr, ). The yellow and red regions in the left and right panels of Fig.11 represent the excluded regions on chargino-LSP and slepton-LSP mass plane. The authors in Ref. Choudhury:2023eje ; Choudhury:2023yfg have extended the similar $4l+\rm E{\!\!\!/}_{T}$ final states in the context of the electroweak sparticle searches at the 14 TeV high-luminosity LHC (HL-LHC) and the proposed high-energy (27 TeV) upgrade of the LHC (HE-LHC). It has been shown that an optimized cut-based analysis will be able to exclude $m_{\widetilde{\chi}_{1}^{\pm}}$ upto 2.18 TeV at the HL-LHC from wino like $\widetilde{\chi}_{1}^{\pm}\widetilde{\chi}_{1}^{\mp}$ and $\widetilde{\chi}_{1}^{\pm}\widetilde{\chi}_{2}^{0}$ productions Choudhury:2023eje . The authors have further explored a multivariate analysis based on an Extreme Gradient BDT algorithm to improve the results further. The projected $2\sigma$ exclusion limit reaches up to 2.37 TeV and 4.0 TeV at the HL-LHC and HE-LHC, respectively, from the ML-based analysis Choudhury:2023eje (the dark violet color represents the $2\sigma$ reach for HL-LHC in Fig the left panel of Fig.11). In another recent work Choudhury:2023yfg , considering the pair and associated production of mass degenerate sleptons, the authors have studied the sensitivity of similar final state at the future LHC. The right panel of Fig.11 shows the discovery and exclusion reach on L-type slepton masses obtained for nonzero $\lambda_{121}$ and/or $\lambda_{122}$ coupling values Choudhury:2023yfg . It is observed that the projected exclusion limits on slepton and sneutrino masses at the HL-LHC (HE-LHC) are $\sim$ 1.85 (3.0) TeV from ML-based analysis. The ML algorithm shows significant improvement over cut-and-count method (see Fig.11). The work in Ref. Bhattacherjee:2023kxw has addressed the search prospect of long-lived particles in the RPV SUSY scenarios from the electroweakino pair production, where the LSP decays via UDD-type couplings. Utilizing the Gradient Boosting algorithm within XGBoost toolkit, the authors obtained that wino-like (higgsino like) $\widetilde{\chi}_{2}^{0}/\widetilde{\chi}_{1}^{\pm}$ with a mass of 1900 (1600) GeV and $\widetilde{\chi}_{1}^{0}$ with a mass greater than 800 (700) GeV can be probed for decay length ranging from 1 cm to 200 cm Bhattacherjee:2023kxw .
@@ 4 Decision Tree algorithms
The decision trees Breiman:1984jka are one of the simplest and most widely used classification and regression tools, which use inverted tree like structure with root on the top, to predict the value of an output by applying binary cuts in the feature space. Breiman et al. breiman ; Breiman:1984jka proposed “Classification and Regression Trees” (CART) algorithm for the implementation of decision trees (DT). The CART algorithm begins with the root-node consisting of the entire training dataset. It then proceeds to split the parent node into several branches or child nodes until the model can make a decision regarding the class of a given data point or it satisfies a predetermined stopping criteria. Through this binary splitting, the algorithm effectively partition the data into subsets. When the splitting criterion is met by the node, it is termed as a leaf.
Age Chest pain Blood pressure weight Diabetic Risk of Heart
(diastolic in mmHg) (Kg) disease
27 No 80 75 No low
65 No 70 71 No low
45 Yes 90 85 Yes high
80 Yes 78 55 Yes high
54 No 100 94 Yes high
40 No 80 68 Yes low
Table 2: A dataset with 6 entries whether a person suffers from heart disease or not. The right most column represents the final output whether a person suffers from heart disease based on the input features.
We demonstrate this process with a medical science example in the Table. 2 regarding the risk of heart disease of a person. In medical science, the risk of being diagnosed with a heart disease primarily depends on the person’s age, blood pressure, weight, whether the person is diabetic or not etc [^9] For the prediction of heart disease using ML algorithms see Ref. hossain2023heart . The initial five columns denote input features, while the last column signifies the resulting outcome. A decision tree, depicted in Fig. 12, is constructed using this dataset. In the tree visualization, the root node is depicted in brown, leaf nodes in green, and general nodes in blue.
@@ 40 No 80 68 Yes low
Figure 12: A decision tree based on the dataset given in Table. 2 is shown here. The root node is marked with brown color. Ordinary nodes and leaf nodes are denoted using blue and green colors, respectively.
@@
In the context of HEP, the goal for DT algorithm is mainly to classify an entry as a signal or background. For classification problem, the journey from the root-node to the leaf-node signifies a series of cuts which determines whether an entry is classified as signal or background depending upon the characteristics of the leaf-node. Each leaf can be assigned a purity value, calculated as $p=S/(S+B)$ where $S$ and $B$ represent the sum of weights of the signal and background events contained within that leaf. Depending upon this value of purity the events in that leaf can be identified as signal or background ( if $p>0.5$, it is identified as signal; otherwise, it is classified as background.) At each node, the feature and its corresponding threshold value determine the subsequent splitting of the node. The selection of the feature and its threshold value depends on the decrease in impurity. The commonly used impurity functions include the misclassification error , the cross entropy breiman ; Breiman:1984jka and the Gini index of diversity gini ; ceriani2012origins , all of which are determined by the purity of signal and background. Among these impurity function, the most popular one is the Gini index which performs similar to entropy. In HEP, generally we try to optimize the signal significance $S/\sqrt{S+B}$ or $S/\sqrt{B}$ by applying Cross section significance or excess significance which is defined as $-S^{2}/(S+B)$ and $-S^{2}/B$ respectively. On the other hand, regression tress trees split nodes based on minimizing the Sum of Squared Errors (SSE). The tree-growing process is stopped using stopping criteria. The stopping condition may include several criteria, such as reaching a maximum tree depth, maximum number of leaf nodes, ensuring a minimum number of instances within each leaf node, insufficient improvement through further splitting, or achieving complete splitting, where all events within the node belong to the same class. Sometimes, even before reaching the maximum allowed depth or number of nodes, early stopping criteria are used to prevent overfitting and improve the generalization.
@@ 4.2 AdaBoost
Decision trees are the base learner for Random Forest and many other algorithms. The learners can be weak or strong depending on their prediction accuracy. Poor prediction accuracy is associated with weak learners, whereas a strong learner yields better model prediction accuracy. While it is hard to predict an output only through a weak learner, a combination of such weak learners can be made into a strong learner. At each step, the misclassified instances are provided with a larger weightage, resulting in better accuracy at the final step. This process is called “Boosting”. AdaBoost freund1999short ; wang2012adaboost works on the basic principle of the boosted decision tree (BDT) algorithm. The term AdaBoost originates from Adaptive Boosting, as the misclassified events are given a larger weightage, while the correctly classified events are given a smaller weightage in this algorithm. Freund and Schapire introduced the AdaBoost algorithm Freund:1997xna in 1995 and from then, it has been widely used in many scenarios including the HEP analysis.
@@
A set of labeled training data $\{x_{i},\hat{y}_{I}\}$ where $x_{i}$ denotes the input features and $\hat{y}_{i}$ stands for the label, is fed into the algorithm. Here $i=1,2,3...m$, and each $x_{i}$ and $\hat{y}_{i}$ belong to domains $X$ and $Y$, respectively ($x_{i}\in X,~{}\hat{y}_{i}\in Y=\{-1,+1\}$). For simplicity, here we have chosen binary classification to elaborate on, and the classes are denoted as $+1$ and $-1$. The base learner is repeatedly called by the algorithm for $t=1,\dots,T$ iteration. It may be noted that $T$ denotes the hyperparameter number of trees/estimators. The weight of the $i^{\text{th}}$ sample in $t^{\text{th}}$ iteration is denoted by $w_{t}(i)$. In the beginning, all samples are assigned the same weight. At each iteration, the learner is compelled to focus on the incorrectly classified event by giving larger weightage to the incorrectly classified events. The learner constructs a weak hypothesis $h_{t}:X\rightarrow\{-1,+1\}$ for $w_{t}$ with the lowest error, where the error is given by $\epsilon_{t}=\sum_{i\colon h_{t}(x_{i})\neq\hat{y}_{i}}w_{t}(i)$. The algorithm next assigns a parameter $\alpha_{t}=\eta\ln(\frac{1-\epsilon_{t}}{\epsilon_{t}})$ to the weak hypothesis, $h_{t}$, which is a measure of performance of $h_{t}$. Here, $\eta$ acts as the learning rate or shrinkage coefficient, which is a measure of the strength of boosting. It is worth noting that for $\epsilon_{t}\leq 1/2$, $\alpha_{t}\geq 0$, and $\alpha_{t}$ increases with decreasing $\epsilon_{t}$. The weight, $w_{t}$, is then modified. A larger weight (multiplying $w_{t}$ by $e^{\alpha_{t}}$) is given for the misclassified events ($h_{t}(x_{i})\neq\hat{y}_{i}$) and a smaller weight (multiplying $w_{t}$ by $e^{-\alpha_{t}}$) is given to the correctly classified events ($h_{t}(x_{i})=\hat{y}_{i}$). After completion of $T^{\text{th}}$ iteration (stopping criterion), the weighted majority of all $T$ weak hypotheses is chosen to be the final hypothesis or strong learner, $H(x)$. The flow chart of the AdaBoost algorithm is presented in Fig. 15.
The ability of AdaBoost to create a strong learner by combining several weak learners by weighted data sampling makes it more efficient than the Random forest algorithm. Furthermore, the AdaBoost training process leads to faster convergence in many scenarios. However, while dealing with complex datasets or when the base learner is too complex, AdaBoost is susceptible to overfitting.
@@ 4.4 LightGBM
Simple gradient-boosting techniques have been quite successful so far. A common attribute of GBDT is to enumerate over all instances in dataset for each feature to find the optimal split. As data complexity and size increase, GBDT faces challenges in efficiency and accuracy. To overcome this issue, novel techniques are implemented upon simple GBDT algorithm, namely, (i) leaf-wise tree growth, (ii) Gradient-based one-sided sampling (GOSS), (iii) histogram splitting, and (iv) Exclusive feature bundling (EFB). The new algorithm that incorporates these new techniques with GBDT, developed by Microsoft, is referred to as Light Gradient Boosting Machine-learning or LightGBM ke2017lightgbm . In the following, we discuss the key concepts of LightGBM:
@@
The flow-chart of LightGBM algorithm is shown in Fig. 18. LightGBM algorithm turns out to be more efficient in terms of time consumption compared with XGBoost while maintaining a competitive accuracy. Due to its histogram-based approach, LightGBM uses less memory, making it more efficient and scalable while dealing with complex large dataset. Nevertheless, as LightGBM splits the trees leaf-wise, it can produce much more complex tree structures, resulting in overfitting in some cases.

### v2

@@ 3.1.2 Searches for RPV SUSY scenarios using BDT
In the presence of non-zero trilinear couplings $\lambda_{ijk}$, the LSP decays into $l_{k}^{\prime\pm}l_{i/j}^{\prime\mp}\nu_{j/i}$, where $l^{\prime}=e,\mu,\tau$. Thus, the final states become leptonically enriched. For different choices of $\lambda_{ijk}$ couplings, the LHC collaborations have already derived limits on chargino and slepton masses (ATLAS:2021yyr, ). The yellow and red regions in the left and right panels of Fig.11 represent the excluded regions on chargino-LSP and slepton-LSP mass plane. The authors in Ref. Choudhury:2023eje ; Choudhury:2023yfg have extended the similar $4l+\rm E{\!\!\!/}_{T}$ final states in the context of the electroweak sparticle searches at the 14 TeV high-luminosity LHC (HL-LHC) and the proposed high-energy (27 TeV) upgrade of the LHC (HE-LHC). It has been shown that an optimized cut-based analysis will be able to exclude $m_{\widetilde{\chi}_{1}^{\pm}}$ upto 2.18 TeV at the HL-LHC from wino like $\widetilde{\chi}_{1}^{\pm}\widetilde{\chi}_{1}^{\mp}$ and $\widetilde{\chi}_{1}^{\pm}\widetilde{\chi}_{2}^{0}$ productions Choudhury:2023eje . The authors have further explored a multivariate analysis based on an Extreme Gradient BDT algorithm to improve the results further. The projected $2\sigma$ exclusion limit reaches up to 2.37 TeV and 4.0 TeV at the HL-LHC and HE-LHC, respectively, from the ML-based analysis Choudhury:2023eje (the dark violet color represents the $2\sigma$ reach for HL-LHC in Fig the left panel of Fig.11). In another recent work Choudhury:2023yfg , considering the pair and associated production of mass degenerate sleptons, the authors have studied the sensitivity of similar final state at the future LHC. The right panel of Fig.11 shows the discovery and exclusion reach on L-type slepton masses obtained for nonzero $\lambda_{121}$ and/or $\lambda_{122}$ coupling values Choudhury:2023yfg . It is observed that the projected exclusion limits on slepton and sneutrino masses at the HL-LHC (HE-LHC) are $\sim$ 1.85 (3.0) TeV from ML-based analysis. The ML algorithm shows significant improvement over the cut-and-count method (see Fig.11). The work in Ref. Bhattacherjee:2023kxw has addressed the searching prospect of long-lived particles in the RPV SUSY scenarios from the electroweakino pair production, where the LSP decays via UDD-type couplings. Utilizing the Gradient Boosting algorithm within XGBoost toolkit, the authors obtained that wino-like (higgsino like) $\widetilde{\chi}_{2}^{0}/\widetilde{\chi}_{1}^{\pm}$ with a mass of 1900 (1600) GeV and $\widetilde{\chi}_{1}^{0}$ with a mass greater than 800 (700) GeV can be probed for decay length ranging from 1 cm to 200 cm Bhattacherjee:2023kxw .
@@ 4 Decision Tree algorithms
The decision trees Breiman:1984jka are one of the simplest and most widely used classification and regression tools, which use an inverted tree like structure with a root on the top, to predict the value of an output by applying binary cuts in the feature space. Breiman et al. breiman ; Breiman:1984jka proposed “Classification and Regression Trees” (CART) algorithm for the implementation of decision trees (DT). The CART algorithm begins with the root-node consisting of the entire training dataset. It then proceeds to split the parent node into several branches or child nodes until the model can make a decision regarding the class of a given data point or it satisfies a predetermined stopping criteria. Through this binary splitting, the algorithm effectively partitions the data into subsets. When the splitting criterion is met by the node, it is termed as a leaf.
We demonstrate this process with an example. Let us consider a gluino pair production channel at the LHC, where each gluino decays into one stable neutralino (LSP) and two quarks ($\tilde{g}\rightarrow q\bar{q}\widetilde{\chi}_{1}^{0}$), resulting in a final state consisting of four jets associated with a large missing transverse energy ($\rm E{\!\!\!/}_{T}$). It should be noted that for the SUSY signal, the final states will consist of jets with large momentum and large values of effective mass ($m_{eff}$) compared to the SM backgrounds.
$p_{T}^{j_{1}}$ GeV $N_{b}$ in GeV $\rm E{\!\!\!/}_{T}$ in GeV $m_{eff}$ in GeV Label
240.2 0 738.1 2112.1 $s$
110.4 1 52.9 352.4 $b$
98.3 0 348.0 1578.2 $s$
87.7 1 68.4 345.7 $b$
61.3 1 27.3 130.6 $b$
129.8 1 102.5 812.1 $b$
Table 2: Dataset to classify signal and background for four input features. The rightmost column (in gray) represents the label, whether it is a signal ($s$) or background ($b$).
@@
Figure 12: A decision tree based on the dataset given in Table. 2 is shown here. The root node is marked with brown color. Ordinary and leaf nodes are denoted using blue and green colors, respectively.
Here, $m_{eff}$ denotes the scalar sum of $p_{T}$ of jets having $p_{T}\geq 50$ GeV added with $\rm E{\!\!\!/}_{T}$. In this case, $t\bar{t}$ + jets act as the most dominant SM background. We label the supersymmetric signal as ‘$s$’ and the SM background as ‘$b$’. To separate signal and background, four variables (input features) are considered, namely transverse momentum of leading jet ($p_{T}^{j_{1}}$), number of b-tagged jets in an event ($N_{b}$), missing transverse energy ($\rm E{\!\!\!/}_{T}$), and, effective mass ($m_{eff}$). A sample dataset is illustrated in Table 2. A decision tree based on Table 2 is shown in Fig. 12. In the tree visualization, the root node is depicted in brown, leaf nodes in green, and ordinary nodes in blue.
@@
In the context of HEP, the goal for DT algorithm is mainly to classify an entry as a signal or background. For the classification problem, the journey from the root-node to the leaf-node signifies a series of cuts that determines whether an entry is classified as signal or background depending upon the characteristics of the leaf-node. Each leaf can be assigned a purity value, calculated as $p=S/(S+B)$ where $S$ and $B$ represent the sum of weights of the signal and background events contained within that leaf. Depending upon this value of purity, the events in that leaf can be identified as signal or background ( if $p>0.5$, it is identified as signal; otherwise, it is classified as background.) At each node, the feature and its corresponding threshold value determine the subsequent splitting of the node. The selection of the feature and its threshold value depends on the decrease in impurity. The commonly used impurity functions include the misclassification error , the cross entropy breiman ; Breiman:1984jka and the Gini index of diversity gini ; ceriani2012origins , all of which are determined by the purity of signal and background. Among these impurity functions, the most popular one is the Gini index, which performs similarly to entropy. In HEP, generally, we try to optimize the signal significance $S/\sqrt{S+B}$ or $S/\sqrt{B}$ by applying Cross section significance or excess significance which is defined as $-S^{2}/(S+B)$ and $-S^{2}/B$ respectively. On the other hand, regression tress trees split nodes based on minimizing the Sum of Squared Errors (SSE). The tree-growing process is stopped using stopping criteria. The stopping condition may include several criteria, such as reaching a maximum tree depth, maximum number of leaf nodes, ensuring a minimum number of instances within each leaf node, insufficient improvement through further splitting, or achieving complete splitting, where all events within the node belong to the same class. Sometimes, even before reaching the maximum allowed depth or number of nodes, early stopping criteria are used to prevent overfitting and improve the generalization.
@@ 4.1 Random Forest
Due to ensemble method and averaging over multiple decision trees, the chances of overfitting get reduced in Random Forest. However, overfitting can still arise due to a very large number of trees, trees with a large number of splits, etc. Fixing the maximum depth of the trees (by setting max_depth parameter) or reducing the number of features considered at each split (max_features) can be useful to prevent overfitting by avoiding too complex model. Another way to do so is by controlling the minimum samples required to split a node (min_samples_leaf) or by applying penalties on large trees. When the feature space is too large, the model becomes complex. Sometimes, few features merely contribute to the final outcome compared to other features in the input feature set. Omitting those less contributing features and only taking the important ones by employing Principle Component Analysis (PCA) can effectively bring down the overfitting issue. To mitigate the effect of noisy data present, one can diversify the training data, hence reducing the overfitting issue.
@@ 4.2 AdaBoost
Decision trees are the base learner for Random Forest and many other algorithms. The learners can be weak or strong depending on their prediction accuracy. Poor prediction accuracy is associated with weak learners, whereas a strong learner yields better model prediction accuracy. While it is hard to predict an output only through a weak learner, a combination of such weak learners can be made into a strong learner. At each step, the misclassified instances are provided with a larger weightage, resulting in better accuracy at the final step. This process is called “Boosting”. AdaBoost freund1999short ; wang2012adaboost works on the basic principle of the boosted decision tree (BDT) algorithm. The term AdaBoost originates from Adaptive Boosting, as the misclassified events are given a larger weightage, while the correctly classified events are given a smaller weightage in this algorithm. Freund and Schapire introduced the AdaBoost algorithm Freund:1997xna in 1995, and from then, it has been widely used in many scenarios, including the HEP analysis.
@@
A set of labeled training data $\{x_{i},\hat{y}_{i}\}$ where $x_{i}$ denotes the input features and $\hat{y}_{i}$ stands for the label, is fed into the algorithm. Here $i=1,2,3...m$, and each $x_{i}$ and $\hat{y}_{i}$ belong to domains $X$ and $Y$, respectively ($x_{i}\in X,~{}\hat{y}_{i}\in Y=\{-1,+1\}$). For simplicity, here we have chosen binary classification to elaborate on, and the classes are denoted as $+1$ and $-1$. The base learner is repeatedly called by the algorithm for $t=1,\dots,T$ iteration. It may be noted that $T$ denotes the hyperparameter number of trees/estimators. The weight of the $i^{\text{th}}$ sample in $t^{\text{th}}$ iteration is denoted by $w_{t}(i)$. In the beginning, all samples are assigned the same weight. At each iteration, the learner is compelled to focus on the incorrectly classified event by giving larger weightage to the incorrectly classified events. The learner constructs a weak hypothesis $h_{t}:X\rightarrow\{-1,+1\}$ for $w_{t}$ with the lowest error, where the error is given by $\epsilon_{t}=\sum_{i\colon h_{t}(x_{i})\neq\hat{y}_{i}}w_{t}(i)$. The algorithm next assigns a parameter $\alpha_{t}=\eta\ln(\frac{1-\epsilon_{t}}{\epsilon_{t}})$ to the weak hypothesis, $h_{t}$, which is a measure of performance of $h_{t}$. Here, $\eta$ acts as the learning rate or shrinkage coefficient, which is a measure of the strength of boosting. It is worth noting that for $\epsilon_{t}\leq 1/2$, $\alpha_{t}\geq 0$, and $\alpha_{t}$ increases with decreasing $\epsilon_{t}$. The weight, $w_{t}$, is then modified. A larger weight (multiplying $w_{t}$ by $e^{\alpha_{t}}$) is given for the misclassified events ($h_{t}(x_{i})\neq\hat{y}_{i}$) and a smaller weight (multiplying $w_{t}$ by $e^{-\alpha_{t}}$) is given to the correctly classified events ($h_{t}(x_{i})=\hat{y}_{i}$). After completion of $T^{\text{th}}$ iteration (stopping criterion), the weighted majority of all $T$ weak hypotheses is chosen to be the final hypothesis or strong learner, $H(x)$. The flow chart of the AdaBoost algorithm is presented in Fig. 15. The ability of AdaBoost to create a strong learner by combining several weak learners by weighted data sampling makes it more efficient than the Random forest algorithm. Furthermore, the AdaBoost training process leads to faster convergence in many scenarios. However, AdaBoost tends to overfit when the data is too noisy and when the weak learners become too complicated. AdaBoost is prone to assigning too much weight to misclassified instances in the high noise regime, producing skewed data distribution. Accommodating regularization techniques into the algorithm can resolve this issue. Similar to Random Forest model, by tuning the hyperparameters, such as max_depth, max_features, can be helpful to reduce overfitting. Also, early stopping, cross-validation, and adjustment of the number of iterations can prevent the algorithm from overfitting.
@@ 4.3 XGBoost
Similar to other decision tree models, XGBoost is also prone to overfitting for a very large number of trees and/or deep trees, small dataset, high learning rate, etc. One can tweak the hyperparameters like max_depth, learning_rate (shrinks the feature weights at each step of boosting), colsample_bytree (to introduce random subsampling of columns), etc. A quantity of the minimum sum of instance weight is assigned for partition. This is denoted as min_child_weight. While splitting a node, if that node comes up with the sum of instances less than the assigned value of min_child_weight, the tree-building process will come to a halt. The $\gamma$ parameter, which is the minimum loss reduction for a partition, can also be changed to reduce model complexity. A higher value for min_child_weight or $\gamma$ makes the algorithm more conservative. One can add randomness to the data by randomly selecting the fraction of the training sample to be fed into the algorithm during training by adjusting subsample parameter. Like AdaBoost, cross-validation and early stopping techniques are also helpful for reducing overfitting.
@@ 4.4 LightGBM
Simple gradient-boosting techniques have been quite successful so far. A common attribute of GBDT is to enumerate all instances in the dataset for each feature to find the optimal split. As data complexity and size increase, GBDT faces challenges in efficiency and accuracy. To overcome this issue, novel techniques are implemented upon simple GBDT algorithm, namely, (i) leaf-wise tree growth, (ii) Gradient-based one-sided sampling (GOSS), (iii) histogram splitting, and (iv) Exclusive feature bundling (EFB). The new algorithm that incorporates these new techniques with GBDT, developed by Microsoft, is referred to as Light Gradient Boosting Machine-learning or LightGBM ke2017lightgbm . In the following, we discuss the key concepts of LightGBM:
@@
The flow-chart of LightGBM algorithm is shown in Fig. 18. LightGBM algorithm turns out to be more efficient in terms of time consumption compared with XGBoost while maintaining a competitive accuracy. Due to its histogram-based approach, LightGBM uses less memory, making it more efficient and scalable while dealing with complex large dataset. Nevertheless, as LightGBM splits the trees leaf-wise, it can produce much more complex tree structures, resulting in overfitting in some cases. Similar to other BDT models, adjusting the parameters like the number of trees, max_depth, learning_rate, subsample is quite useful to prevent overfitting. Controlling the number of bins in which feature values are bucketed can reduce the chances of overfitting. This can be done by setting the max_bin parameter properly. To avoid complex tree structures, the minimum data in one leaf and the sum of hessian in one leaf are useful, along with the number of leaves in one tree. This procedure can be done using min_data_in_leaf, min_sum_hessian_in_leaf, and num_leaves parameters, respectively. While partitioning, the minimum gain to perform a split can be done using min_gain_to_split. In addition, implementing techniques such as regularization and cross-validation also aid in preventing overfitting.